A plan of a shopping complex is drawn to a scale of 1 cm to 5 m. Express the scale of the plan in the form 1: n . Find the perimeter and area of a rectangular shop space which measures 2.4cm by 4cm on the plan.
Answer by Alan3354(69443)

A plan of a shopping complex is drawn to a scale of 1 cm to 5 m. Express the scale of the plan in the form 1: n . Find the perimeter and area of a rectangular shop space which measures 2.4cm by 4cm on the plan.
-----------
i cm to 5 m = 1 cm to 500 cm
= 1:500
-----------
The perimeter = 2w+2l = 2(2.4 + 4)
p = 6.4 cm
at 1:500 that's 3200 cm
------------
The area is (2.4*500)*(4*500)
A = 1200*2000
A = 2400000 sq cms
A = 240 sq meters